OpenStudy (anonymous):

Algebra question: 3(x+h)^2

OpenStudy (anonymous):

not sure what to do first

OpenStudy (anonymous):

3(x+h)(x+h) 3(x^2 + 2xh + h^2) 3x^2 + 6xh + 3h^2 if your just expanding the brackets, is that what your after?

OpenStudy (anonymous):

uhh, well I am doing definition of derivative and i'm not sure I was expecting polynomials

OpenStudy (anonymous):

is that the absolute way of solving that? @JamesWolf

OpenStudy (anonymous):

never mind, thanks

OpenStudy (anonymous):

well what is it equal too? you have 2 unknowns and 1 equation. what are you solving it for. your trying to take the derivative with respect to x? is h constant? sortof need to know these things to help

OpenStudy (anonymous):

derivative will be u = x +h f(x) = 3(u)^2 df(x)/dx = df(x)/du x du/dx = 6(u) x 1 = 6(u) = 6(x+h)

OpenStudy (anonymous):

Thank you, I got it from there :)